An insurance company claims that for x thousand policies, its monthly revenue in dollars is given by R(c)=125x and its monthly cost in dollars is given by C(x) = 100x + 5000

1 Find the break-even point

2 Graph the revenue and cost equations on the same axes

3 From the graph, estimate the revenue and cost when x = 100 (100,000 policies)
Answer #1

C(x) = 100x + 5000

R(c)=125x

a)

break even point is

when

R(C) = C(X)

125x = 100x + 5000

x = 200

break even point = 200000 policies

b)

40000 35000 30000 25000 20000 - 15000 10000 100 140 180 220 260 300

red line is for R(C) and blue line is for C(X)

c) revenue = 125x = 125*100000 =$ 12500000

cost = 100x + 5000 = 100*100000 + 500 =$ 10005000

please revert for doubt...